• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name ZAFAR AHMED

    Log entry time 23:59:21 on November24,2009

    Entry number 302373

    keyword=Swing shift summary

    16:00 Shift startts with Moller measurements.
    16:30 We were not able to see the asymmetry in Moller so asked for controlled access.
    17:00 Everbody is out of the hall but shunt power supply on one of the ARC magnets. Its not responding to remote restet MCC going for local reset if it not wrks they will replace it.
    17:30 MCC is not able to fix shunt module so they called the expert. It will take one hour or may be more.
    20:00 Beam is back.
    23:00 Sasha finished the moller and MCC is doing the HARP scan.
    23:15 HARP scandone. 1H3A = 203/151 and 3H3B = 230/127.
    23:15 Checking the raster with T1 trigger only on left arm. Target is BeO.
    23:30 Raster is good 2.5x3. Now we are moving to LD2 target and start taking production runs.